Help on this question of Linear Algebra, thanks.

Let A= [ 0.7 0.1 0.3 0.9 Find P and D such that A = PDP-1, where D is a diagonal matrix.
Answer #1

Date. Page 1 007 o.1 0.3 A = 1/10 9 3 10 characteristic equation is d2 - ( 16 dt 60 Lo 10 1 0 0 od +60 10 16 22 o To dld-(10page2 3/10 - 7/10 9/10 -1/10 y lo - 3 x 16 7. 2 - 16 Ni= 1/3 1 d = 315 A-3/7) x = 0 Ilse S/10 fo] I 3710 3110 V2 = 1 1 chosepade 3 P = - 113 1 1 P-1 AP 1 In decimal Form P 10.33333 1 I 1 1 o 315 In decimal Form D Ć 1